Hi, we just started learning about quadratic equations today, and I found some problems that I needed help with.
Define a variable and write a quadratic function to describe each situation.
30. the product of two numbers whose sum is 45
31. the area of a circle in terms of its radius
32. the product of two numbers whose difference is 15
any help would be appreciated, thank you!

Answer by Alan3354(69443)

Define a variable and write a quadratic function to describe each situation.
30. the product of two numbers whose sum is 45
x*(45-x) = Product
45x - x^2 = Product
31. the area of a circle in terms of its radius
Area = pi*r^2
32. the product of two numbers whose difference is 15
Product = n*(n+15) = n^2 + 15n